OpenStudy (arbyj):

how do i find the solutions to 7-12x^2 = -1?

OpenStudy (agent47):

Move all the constant terms to one side: -12x^2 = -1 - 7 -12x^2 = -8 Divide everything by -12 x^2 = 8/12 Simplify and take the sqrt

OpenStudy (arbyj):

so would it simplify to (2x-3)(2x+3) ?? @agent47

OpenStudy (freckles):

did you post the right question? I'm not seeing how your new question has anything to do with your old question?

OpenStudy (greatlife44):

Yeah, I think the way @agent47 did it is the best way

OpenStudy (arbyj):

@freckles what do you mean

OpenStudy (freckles):

\[7-12x^2 = -1 \text{ is the equation you posted } \\ \text{ add 1 on both sides } \\ 1+7-12x^2=-1+1 \\ 8-12x^2=0 \\ \text{ divide both sides by } 4 \\ 2-3x^2=0 \\ \text{ you see } 2-3x^2 \text{ doesn't factor to what you have } \\ \text{ so I thought you were asking a different question } \\ \text{ but wasn't totally sure }\]